# Towards reconstruction of finite tensor categories

### v1

@@ Abstract.
We take a first step towards a reconstruction of finite tensor categories using finitely many $F$-matrices. The goal is to reconstruct a finite tensor category from its projective ideal. Here we set up the framework for an important concrete example—the $8$-dimensional Nicholas Hopf algebra $K_{2}$. Of particular importance is to determine its Green ring and tensor ideals. The Hopf algebra $K_{2}$ allows the recovery of $(2+1)$-dimensional Seiberg-Witten TQFT from Hennings TQFT based on $K_{2}$. This powerful result convinced us that it is interesting to study the Green ring of $K_{2}$ and its tensor ideals in more detail. Our results clearly illustrate the difficulties arisen from the proliferation of non-projective indecomposable objects in finite tensor categories.
@@ 1. Introduction
A salient feature of non-semisimple Hopf algebras is the proliferation of indecomposable modules that are not projective. As we will see in Sec. 5 that for the $8$-dimensional Nicholas Hopf algebra $K_{2}$, its indecomposable modules include a continuous family parameterized by the $2$-sphere. Therefore, it is not even clear whether or not that finite tensor categories can be presented by a finite set of matrices as in the case of fusion categories. We prove that the full subcategory of projectives is a complete invariant of a finite tensor category. Hence to present a finite tensor category, it suffices to do so for its projective subcategory. This in principle reduces the presentation, invariants, classification, and open problems of finite tensor categories to these of their projective subcategories. Note that the projective subcategory of a fusion category is itself.
@@
One preparation work for the example $K_{2}$ is a detailed study of its Green ring and tensor ideals. The $8$-dimensional Nicholas Hopf algebra $K_{2}$ allows the recovery of $(2+1)$-dimensional Seiberg-Witten TQFT from Hennings TQFT based on $K_{2}$ [8]. This powerful result convinced us to study the Green ring of $K_{2}$ in more detail. One major goal is to determine the tensor ideals of the Green ring $r(K_{2})$. These results clearly illustrate the difficulties arisen from the proliferation of non-projective indecomposable objects in finite tensor categories. In the future, we plan to classify all finite tensor categories with the same projective fusion rules.
@@ 1.1. Notation and terminology
    $\overline{S}^{\textrm{AC}}$=the additive subcategory whose objects are finite direct sums of elements of $S$ where $S$ is a subset of the FTC $\mathcal{C}$ known as the additive completion of $S$
@@ 2.1.1. Green rings $r(\mathcal{C})$ and projective fusion rings $K_{0}(\mathcal{C})$
The Green rings of the Taft algebras are determined in [2]. Since the Sweedler Hopf algebra $K_{1}$ is a Taft algebra, it follows that the Green ring $r(K_{1})$ of the Sweeder Hopf algebra is $r(K_{1})=\mathbb{Z}[x,y]/I$, where $I$ is the ideal generated by $x^{2}-1$ and $y^{2}-x-xy$ [2]. The Green ring of the Drinfeld double of the Sweedler Hopf algebra is obtained in [3]. The Green rings of $K_{m}$ for $m\geq 2$ is unknown. One of our results is the determination of the Green ring $r(K_{2})$ of $K_{2}$. We deduce this result from a quotient map from $DK_{1}$ to $K_{2}$ and the Green ring of the Drinfeld double of $K_{1}$ [3].
@@ 2.1.3. Projective Auslander algebras
Given an FTC $\mathcal{C}$ and a full list $\{P_{i}\}_{i\in I}$ of its projective indecomposable module representatives, the projective Auslander algebra $A(\mathcal{C})$ is the algebra $A(\mathcal{C})=\operatorname{Hom}(P,P),$ where $P=\oplus_{i\in I}P_{i}$ with the composition of morphisms as the algebra multiplication.
@@ 2.2.3. Projective Auslander algebras of $K_{m}$
Let $P_{0}$ and $P_{1}$ denote $P(0)$ and $P(1)$ respectively for $K_{m}$. Let $A$ denote the Auslander algebra: $A=\operatorname{End}_{K_{n}}(P_{0}\oplus P_{1})$. Let $e_{0}=\frac{1+K}{2}$ and $e_{1}=\frac{1-K}{2}$. Let $W$ denote the set of elements of $K_{m}$: $\xi_{i_{1}}\xi_{i_{2}}\cdots\xi_{i_{k}}$ with $1\leq i_{1}<i_{2}\cdots<i_{k}$ for $0\leq k\leq m$. Then $We_{i}$ forms a basis for $P_{i}$. Then, note that for $r,s\in\{0,1\}$, as vector spaces,
$\hom_{K_{n}}(P_{r},P_{s})$ $\simeq\hom_{K_{n}}(P_{r}\otimes P_{s}^{*},1)$
$\hom_{K_{n}}\simeq(2^{m-1}P_{0}\oplus 2^{m-1}P_{1},1)$
$\hom_{K_{n}}\simeq(2^{m-1}P_{0},1),$
which is dimension $2^{m-1}$. Further, note that $\phi\in\hom_{K_{n}}(P_{r},P_{s})$ is entirely determined by the image of $e_{r}$ since $\phi(we_{r})=w\cdot\phi(e_{r})$. Further, note that for a map $\phi:P_{r}\to P_{s}$,
@@
$=w\cdot(e_{r+|w|})e_{s},$
then for $r=s$, the odd length terms must be zero and for $r=s+1$, the even length terms are zero. Thus, there are $\sum\limits_{i=1}^{\lfloor\frac{m}{2}\rfloor}\begin{pmatrix}m\\ 2i\end{pmatrix}=2^{m-1}$ possible non-zero coefficients, so each must be possible by the dimension of the space of algebra homomorphisms. Further, for $w\in W$, let $\phi^{w}_{r}:P_{r}\to P_{r+|w|}$ denote the map $e_{r}\mapsto we_{r+|w|}$ and let $\phi_{r}^{i}:P(r)\to P(1+r)$ where $e_{r}\mapsto\xi_{i}e_{1+r}$. First, note that the set $\{\phi^{w}_{r}\}$ generate all possible maps from $P_{r}\to P_{s}$, for fixed $r$ and either choice of $s$, so the Auslander algebra is generated by the set $\{\phi_{r}^{w}\}$ with $r\in R$ and $w\in W$. However, for non-empty word, note each $\phi_{r}^{w}$ is generated as a product
$\phi_{r}^{w}=\phi_{r+|w|}^{w_{1}}\cdots\circ\phi_{r+1}^{w_{|w|-1}}\circ\phi_{r}^{w_{|w|}}$
@@ 3.1.1. Full projective subcategory of an abelian category with enough projectives
Suppose $\mathcal{C}$ is an abelian category with enough projectives and possibly some additional structures, and let $\mathcal{P}_{\mathcal{C}}$ denote its full subcategory of projectives. Let $\mathcal{B}_{R}$ be the category with objects additive categories with the same structure and morphisms natural equivalence classes of right exact functors preserving this structure, and let $\mathcal{A}$ denote the full subcategory of $\mathcal{B}_{R}$ with abelian objects. For the case of finite tensor categories, this corresponds to the category with objects locally-finite $\mathbb{C}$-linear rigid monoidal additive categories with bilinear tensor product with maps natural equivalence classes of right exact additive tensor functors and the full subcategory of finite tensor categories.
@@ 3.3. Categorical reconstruction from projectives
If $\mathcal{P}$ is a full projective subcategory of a finite tensor category, then for a map $\phi:P_{i}\to P_{k}$ between projective indecomposables, im $\text{im\,}\phi$ is a simple object if and only if for every map $\psi:P_{l}\to P_{k}$ and $j:P_{k}\to P$ such that $j\circ\psi=0$, then $j\circ\phi=0$ as well where $P_{i},P_{k},P_{l}$ are projective indecomposables and $P$ is projective.
@@ 4.1.1. The map $\pi_{*}:r(K_{2n})\to r(DK_{n})_{0}$
Thus, since $K-\overline{K}$ is in the annihilator of all simple tensors, it is in the annihilator of the entire tensor product. Thus, $[M_{1}],[M_{2}]\in r(DK_{n})_{0}$, then $[M_{1}\otimes M_{2}]\in r(DK_{n})_{0}$. ∎
@@ 5.1.3. Indecomposable modules of $K_{2}$
We separate the indecomposable modules of $K_{2}$ into three types: the simple-projective $\{V(r),P(r)\}$, the Syzygy type $\{\Omega^{s}V(r),\Omega^{-s}V(r)\}$, and the continuous M-type $\{M_{n}(r,\eta)\}$. We deduce the following from [3].
@@
    $M_{s}(r,\eta)\otimes M_{n}(r^{\prime},\eta^{\prime})=\oplus snP(r+r^{\prime})$ if $\eta\neq\eta^{\prime}$, $M_{s}(r,\eta)\otimes M_{n}(r^{\prime},\eta)=\oplus s(n-1)P(r+r^{\prime})\oplus M_{s}(0,\eta)\oplus M_{s}(1,\eta)$
@@ 5.2. Tensor ideals of the Representation Category of $K_{2}$
Recall for any projective $P\in\operatorname{Rep}(K_{2})$ and indecomposable $A\in\operatorname{Rep}(K_{2})$, $A\otimes P$ is projective. Further, note that by Propositions 2.4 ($V(r^{\prime})$), 2.25 and 2.33 $(M_{n}(r^{\prime},\eta^{\prime})$), Corollary 2.22 $(\Omega^{n}(V(r))$, Corollary 2.24 $(\Omega^{-n}(V(r))$ and Corollary 2.10 ($P(r^{\prime})$) of [3], for any indecomposable $A\in\operatorname{Rep}(K_{2})$,
@@
A tensor subcategory $\mathcal{T}$ of $\operatorname{Rep}(K_{2})$ is fully classified by the set $A_{\mathcal{T}}$ where $A_{\mathcal{T}}$ is a subset of the set of all finite subsets of +×$\mathbb{Z}_{+}\times\mathbb{C}P^{1}\times\mathbb{Z}_{+}\times\{0,1\}$ where $T\in A_{\mathcal{T}}$ if and only if $\sum\limits_{(n,\eta,k,r)\in T}kM_{n}(r,\eta)\in\mathcal{T}$.
@@ 5.2.1. Quasi-dominated subcategory of $r(K_{2})$
5.3. Multiplicity modules
Given an FTC $\mathcal{C}$ with its projective ideal $\mathcal{P}$. Let $I_{\mathcal{C}}$ be the index set of isomorphism classes of indecomposable projective objects of $\mathcal{C}$, and $\{P_{i}\}_{i\in I_{\mathcal{C}}}$ a complete set of representatives of indecomposable projective objects of $\mathcal{C}$.
Suppose $I=i_{1}i_{2}\cdots i_{n},i_{j}\in I_{\mathcal{C}}$ is a multi-index, then we denote $P_{I}=P_{i_{1}}\otimes P_{i_{2}}\otimes\cdots P_{i_{n}}$ and $|I|=n$. We will denote the projective Auslander algebra as $A=\operatorname{Hom}(P,P)$, where $P=\oplus_{i\in I_{\mathcal{C}}}P_{i}$. For each $i$, the multiplicity module $\operatorname{Hom}(P,P_{i})$ is an $A$-module by precomposition.
5.3.1. Dimensions of multiplicity modules
The projective fusion coefficients $N_{ij}^{k}$ is defined by the projective fusion rules: $P_{i}\otimes P_{j}=\oplus_{k\in I_{\mathcal{C}}}N_{ij}^{k}P_{k}.$
For fusion categories, the fusion coefficient $N_{ij}^{k}=\dim\operatorname{Hom}(P_{i}\otimes P_{j},P_{k})$. But this does not hold anymore for general FTCs. The dimensions of multiplicity modules for three projectives could be different from the fusion coefficients.
Let $M_{i}^{j}$ be the multiplicity defined by $M_{i}^{j}=\operatorname{Hom}(P_{i},P_{j})$. For fusion categories, $M_{i}^{j}$ is the Kronecker delta $\delta_{ij}$. But for FTCs, they can be very general. It follows that
$\dim\operatorname{Hom}(P_{i}\otimes P_{j},P_{k})=\sum_{l\in I_{\mathcal{C}}}M^{k}_{l}N_{ij}^{l}.$
Therefore, the computation of the multiplicities $M_{I}^{J}=\dim\operatorname{Hom}(P_{I},P_{J})$ is more complicated. It would be interesting to find generalized Verlinder-type formulas for the multiplicities.
5.3.2. Properties of the intertwining maps
Recall the simple and projective modules $V(r)=V(1,r)$, $V(2,r)$, and $P(r)$ of $DK_{1}$, where $r\in\mathbb{Z}_{2}$ are explicitly given as follows. We first identify interesting maps between these $6$ modules.
The module $V(1,r)$ has a basis $\{v\}$ and has a module structure
$\begin{array}[]{cccc}a.v=0&d.v=0&b.v=(-1)^{r}v&c.v=(-1)^{r}v\\ \end{array}$
The two two-dimensional modules $V(2,r)$ have a basis $\{v_{1},v_{2}\}$ such that
$\begin{array}[]{llll}a.v_{1}=v_{2}&d.v_{1}=0&b.v_{1}=(-1)^{r}v_{1}&c.v_{1}=(-1)^{r+1}v_{1}\\ a.v_{2}=0&d.v_{2}=2v_{1}&b.v_{2}=(-1)^{r+1}v_{2}&c.v_{2}=(-1)^{r}v_{2}.\end{array}$
Finally, the projective covers of $V(1,r)$ are $P(r)$, which have a basis of $\{v_{1},v_{2},v_{3},v_{4}\}$ and module structure given by
$\begin{array}[]{llll}a.v_{1}=v_{2}&d.v_{1}=v_{3}&b.v_{1}=(-1)^{r}v_{1}&c.v_{1}=(-1)^{r}v_{1}\\ a.v_{2}=0&d.v_{2}=-v_{4}&b.v_{2}=(-1)^{r+1}v_{2}&c.v_{2}=(-1)^{r+1}v_{2}\\ a.v_{3}=v_{4}&d.v_{3}=0&b.v_{3}=(-1)^{r+1}v_{3}&c.v_{3}=(-1)^{r+1}v_{3}\\ a.v_{4}=0&d.v_{4}=0&b.v_{4}=(-1)^{r}v_{4}&c.v_{4}=(-1)^{r}v_{4}\\ \end{array}$
First, consider maps between these six modules. Let $\phi:V(r)\to P(r^{\prime})$. Then,
$\phi(v)=\lambda_{1}v_{1}+\lambda_{2}v_{2}+\lambda_{3}v_{3}+\lambda_{4}v_{4}.$
Since $\phi(a.v)=0$, then $a.\phi(v)=0$ as well, so $\lambda_{1}v_{2}+\lambda_{3}v_{4}=0$. Thus, $\lambda_{1}=\lambda_{3}=0$. Also, $d.v=0$, so $d.\phi(v)=0$, which implies that $\lambda_{2}=0$ as well. Finally, since $b.v=(-1)^{r}v$, then $b.\phi(v)=(-1)^{r}\lambda_{1}v_{1}$, which implies that $\lambda_{1}=0$ if $r^{\prime}\neq r$.
Next, consider a map $\phi:P(r)\to V(r^{\prime})$. Then suppose $v_{1}\mapsto cv$. Then, since $a.v_{1}=v_{2}$ and $a.v=0$, then $\phi(v_{2})=0$. Similarly, since $d.v_{1}=v_{3}$ and $d.v=0$, then $\phi(v_{3})=0$. Since $\phi(v_{3})=0$ and $a.v_{3}=v_{4}$, then $\phi(v_{4})=0$ as well. Finally, since $b.v_{1}=(-1)^{r}v_{1}$, then $b.\phi(v_{1})=(-1)^{r}cv$ which implies $c=0$ if $r\neq r^{\prime}$.
Consider a map $\phi:P(r)\to V(2,r^{\prime})$. Then, since $b.v=c.v$ for all $v\in P(r)$, then $b.\phi(v)=c.\phi(v)$ for all $v\in P(r)$, but the only possible $\phi(v)$ that satisfies this is precisely zero. By the same argument, any map from $V(r)\to V(2,r^{\prime})$ is the zero map as well.
If we consider a map from $\phi:V(2,r)\to P(r^{\prime})$, then since $b.v=(-1)c.v$ for all $v\in V(2,r)$, then $b.\phi(v)=-c.\phi(v)$, and the only possible element $\phi(v)$ that satisfies this property is zero. By the same argument, any map from $V(2,r)\to V(r^{\prime})$ is the zero map as well.
Now consider a map $\phi:P(r)\to P(r+1)$. Suppose
$\phi(v_{1})=\lambda_{1}v_{1}+\lambda_{2}v_{2}+\lambda_{3}v_{3}+\lambda_{4}v_{4}.$
Then, $b.\phi(v_{1})=(-1)^{r}\phi(v_{1})$, but
$(-1)^{r}\phi(v_{1})$ $=b.\phi(v_{1})$
$(-1)^{r}\phi(v_{1})$ $=b.(\lambda_{1}v_{1}+\lambda_{2}v_{2}+\lambda_{3}v_{3}+\lambda_{4}v_{4})$
$(-1)^{r}(\lambda_{1}v_{1}+\lambda_{2}v_{2}+\lambda_{3}v_{3}+\lambda_{4}v_{4})$ $=(-1)^{r+1}\lambda_{1}v_{1}+(-1)^{r}\lambda_{2}v_{2}+(-1)^{r}\lambda_{3}v_{3}+(-1)^{r+1}\lambda_{4}v_{4}$
so $\lambda_{1}=\lambda_{4}=0$. Then, $\phi(v_{2})=\phi(a.v_{1})$, so
$\phi(v_{2})=\lambda_{3}v_{4}$
and $\phi(v_{3})=\phi(d.v_{1})$, so
$\phi(v_{3})=-\lambda_{2}v_{4}.$
Finally, $\phi(v_{4})=\phi(a.v_{3})=0$.
Lastly, consider a map $\phi:P(r)\to P(r)$. Suppose
$\phi(v_{1})=\lambda_{1}v_{1}+\lambda_{2}v_{2}+\lambda_{3}v_{3}+\lambda_{4}v_{4}.$
By the same argument as before, since $b.\phi(v_{1})=(-1)^{r}\phi(v_{1})$, then $\lambda_{2}=\lambda_{3}=0$. Since $v_{2}=a.v_{1}$, then $\phi(v_{2})=\lambda_{1}v_{2}$. Since $v_{3}=d.v_{1}$, then $\phi(v_{3})=\lambda_{1}v_{3}$. Finally, since $a.v_{3}=v_{4}$, then $\phi(v_{4})=\lambda_{1}v_{4}$.
5.3.3. Fusion spaces $\operatorname{Hom}(P_{i}\otimes P_{j},P_{k})$ for the Nicholas Hopf algebras $K_{m}$
In this subsection, we obtain a classification of maps $P_{1}\otimes P_{2}\to P_{3}$, where $P_{i}\in\{P(0),P(1)\}$ for Nicholas Hopf algebras $K_{m}$. These maps will be used in the future to skeletalize the projective sub-categories.
First, note that $P_{1}\otimes P_{2}\cong 2^{m-1}P(0)\oplus 2^{m-1}P(1)$ by [Proposition 7.2.1] [1]. Thus,
$\hom_{K_{m}}(P_{1}\otimes P_{2},P_{3})$ $\cong\hom_{K_{m}}(2^{m-1}P(0)\oplus 2^{m-1}P(1),P_{3})$
$\cong\bigoplus_{j=1}^{2^{m-1}}\hom_{K_{m}}(P(0),P_{3})\oplus\bigoplus_{j=1}^{2^{m-1}}\hom_{K_{m}}(P(1),P_{3})$
$\cong\bigoplus_{j=1}^{2^{m-1}}\hom_{K_{m}}(P(0)\otimes P_{3}^{*},V(0))\oplus\bigoplus_{j=1}^{2^{m-1}}\hom_{K_{m}}(P(0)\otimes P_{3}^{*},V(0))$
$\cong\bigoplus_{j=1}^{2^{m}}\hom_{K_{m}}(2^{m-1}P(0)\oplus 2^{m-1}P(1),V(0))$
$\cong\bigoplus_{j=1}^{2^{2m-1}}\hom_{K_{m}}(P(0),V(0)),$
so it is of dimension $2^{2m-1}$.
Let $e_{i}=\frac{(1+K)}{2}$ if $P_{i}=P(0)$ and $e_{i}=\frac{(1-K)}{2}$ if $P_{i}=P(1)$. Now, using the notation of [1], let $c_{\omega,\omega^{\prime}}=f(\omega e_{1}\otimes\omega^{\prime}e_{2})$ for $\omega,\omega^{\prime}\in W$. Then, since $K\cdot(\omega e_{1}\otimes\omega^{\prime}e_{2})=\pm\omega e_{1}\otimes\omega^{\prime}e_{2}$, then $f(\omega e_{1}\otimes\omega^{\prime}e_{2})$ must be in the subspace of $P_{3}$ of dimension $2^{m-2}$ where $K\cdot v$ is the corresponding ±$\pm v$. Also, note that
$\xi_{i}\cdot c_{\omega,\omega^{\prime}}$ $=\xi_{i}\cdot f(\omega e_{1}\otimes\omega^{\prime}e_{2})$
$=\xi_{i}\cdot f(\omega e_{1}\otimes\omega^{\prime}e_{2})$
$=f(K\omega e_{1}\otimes\xi_{i}\omega^{\prime}e_{2}+\xi_{i}\omega e_{1}\otimes\omega^{\prime}e_{2})$
=±$=\pm c_{\omega,\xi_{i}\omega^{\prime}}\pm c_{\xi_{i}\omega,\omega^{\prime}}.$
Therefore, $c_{\xi_{i}\omega,\omega^{\prime}}$ is entirely determined by $c_{\omega,\omega^{\prime}}$ and $c_{\omega,\xi\omega^{\prime}}$. Additionally, if $\xi_{i}\in\omega^{\prime}$, then $\xi_{i}\omega^{\prime}=0$, so
$\xi_{i}\cdot c_{\omega,\omega^{\prime}}=\pm c_{\xi_{i}\omega,\omega^{\prime}}.$
Thus, if $\xi_{i}\in\omega^{\prime}$, then $c_{\xi\omega,\omega^{\prime}}$ is entirely determined by $c_{\omega,\omega^{\prime}}$. Thus the set $c_{1,\omega}$ determine the entire map. This is a set of $2^{m}$ elements and each element lives in a determined subspace of dimension $2^{m-1}$. Thus, the solution space to this set of equations is dimension $2^{2m-1}$, and since the dimension of the space of homomorphisms is also $2^{2m-1}$, then this must be the entire space. Thus, the homomorphism is uniquely determined by the $2^{m}$ $c_{1,\omega}$ images and each of these images is in the corresponding $2^{m-1}$ dimensional subspace of $P_{3}$ such that the action of $K$ agrees on both sides.
These maps will serve as basis elements for the multiplicity spaces and play an important role for the pentagon equations of the projective sub-category of $r(K_{2})$. It would be interesting to solve these pentagons and decide whether or not that the projective Auslander algebra and the projective fusion rules of $K_{2}$ uniquely determine $\operatorname{Rep}(K_{2})$ as an FTC.

### v2

@@ Abstract.
We take a first step towards a reconstruction of finite tensor categories using finitely many $F$-matrices. The goal is to reconstruct a finite tensor category from its projective ideal. Here we set up the framework for an important concrete example—the $8$-dimensional Nicholas Hopf algebra $K_{2}$. Of particular importance is to determine its Green ring and tensor ideals. The Hopf algebra $K_{2}$ allows the recovery of $(2+1)$-dimensional Seiberg-Witten TQFT from Hennings TQFT based on $K_{2}$. This powerful result convinced us that it is interesting to study the Green ring of $K_{2}$ and its tensor ideals in more detail. Our results clearly illustrate the difficulties arisen from the proliferation of non-projective reducible indecomposable objects in finite tensor categories.
@@ 1. Introduction
A salient feature of non-semisimple Hopf algebras is the proliferation of reducible indecomposable modules that are not projective. As we will see in Sec. 5 that for the $8$-dimensional Nicholas Hopf algebra $K_{2}$, its indecomposable modules (up to isomorphisms) include a continuous family parameterized by the $2$-sphere. Therefore, it is not even clear whether or not that finite tensor categories can be presented by a finite set of matrices as in the case of fusion categories. We prove that the full subcategory of projectives is a complete invariant of a finite tensor category. Hence to present a finite tensor category, it suffices to do so for its projective subcategory. This in principle reduces the presentation, invariants, classification, and open problems of finite tensor categories to these of their projective subcategories. Note that the projective subcategory of a fusion category is itself.
@@
One preparation work for the example $K_{2}$ is a detailed study of its Green ring and tensor ideals. The $8$-dimensional Nicholas Hopf algebra $K_{2}$ allows the recovery of $(2+1)$-dimensional Seiberg-Witten TQFT from Hennings TQFT based on $K_{2}$ [8]. This powerful result convinced us to study the Green ring of $K_{2}$ in more detail. One major goal is to determine the tensor ideals of the Green ring $r(K_{2})$. These results clearly illustrate the difficulties arisen from the proliferation of non-projective reducible indecomposable objects in finite tensor categories. In the future, we plan to classify all finite tensor categories with the same projective fusion rules.
@@ 1.1. Notation and terminology
    $\overline{S}^{\textrm{AC}}$=the additive subcategory whose objects are finite direct sums of elements of $S$ where $S$ is a subset of the FTC $\mathcal{C}$ known as the finite additive completion of $S$
@@ 2.1.1. Green rings $r(\mathcal{C})$ and projective fusion rings $K_{0}(\mathcal{C})$
The Green rings of the Taft algebras are determined in [2]. Since the Sweedler Hopf algebra $K_{1}$ is a Taft algebra, it follows that the Green ring $r(K_{1})$ of the Sweeder Hopf algebra is $r(K_{1})=\mathbb{Z}[x,y]/I$, where $I$ is the ideal generated by $x^{2}-1$ and $y^{2}-y-xy$ [2]. The Green ring of the Drinfeld double of the Sweedler Hopf algebra is obtained in [3]. The Green rings of $K_{m}$ for $m\geq 2$ is unknown. One of our results is the determination of the Green ring $r(K_{2})$ of $K_{2}$. We deduce this result from a quotient map from $DK_{1}$ to $K_{2}$ and the Green ring of the Drinfeld double of $K_{1}$ [3].
@@ 2.1.3. Projective Auslander algebras
Given an FTC $\mathcal{C}$ and a full list $\{P_{i}\}_{i\in I}$ of its projective indecomposable module representatives, we define the projective Auslander algebra $A_{\mathcal{C}}$ to be the algebra $A_{\mathcal{C}}=\operatorname{Hom}(P,P),$ where $P=\oplus_{i\in I}P_{i}$ with the composition of morphisms as the algebra multiplication.
@@ 2.2.3. Projective Auslander algebras of $K_{m}$
Let $P_{0}$ and $P_{1}$ denote $P(0)$ and $P(1)$ respectively for $K_{m}$. Let $A$ denote the Auslander algebra: $A=\operatorname{End}_{K_{m}}(P_{0}\oplus P_{1})$. Let $e_{0}=\frac{1+K}{2}$ and $e_{1}=\frac{1-K}{2}$. Let $W$ denote the set of elements of $K_{m}$: $\xi_{i_{1}}\xi_{i_{2}}\cdots\xi_{i_{k}}$ with $1\leq i_{1}<i_{2}\cdots<i_{k}$ for $0\leq k\leq m$. Then $We_{i}$ forms a basis for $P_{i}$. Let $p(n)$ denote the parity function $p(n)=n$ mod $2$. Then, note that for $r,s\in\{0,1\}$, as vector spaces,
$\hom_{K_{m}}(P_{r},P_{s})$ $\simeq\hom_{K_{m}}(P_{r}\otimes P_{s}^{*},1)$
$\simeq\hom_{K_{m}}(2^{m-1}P_{0}\oplus 2^{m-1}P_{1},1)$
$\simeq\hom_{K_{m}}(2^{m-1}P_{0},1),$
which is dimension $2^{m-1}$. Further, note that $\phi\in\hom_{K_{m}}(P_{r},P_{s})$ is entirely determined by the image of $e_{r}$ since $\phi(we_{r})=w\cdot\phi(e_{r})$. Further, note that for a map $\phi:P_{r}\to P_{s}$,
@@
$=w\cdot(e_{r+p(|w|)})e_{s},$
then for $r=s$, the odd length terms must be zero and for $r=s+1$, the even length terms are zero. Thus, there are $\sum\limits_{i=1}^{\lfloor\frac{m}{2}\rfloor}\begin{pmatrix}m\\ 2i\end{pmatrix}=2^{m-1}$ possible non-zero coefficients, so each must be possible by the dimension of the space of algebra homomorphisms. Further, for $w\in W$, let $\phi^{w}_{r}:P_{r}\to P_{r+p(|w|)}$ denote the map $e_{r}\mapsto we_{r+p(|w|)}$ and let $\phi_{r}^{i}:P(r)\to P(1+r)$ where $e_{r}\mapsto\xi_{i}e_{1+r}$. First, note that the set $\{\phi^{w}_{r}\}$ generate all possible maps from $P_{r}\to P_{s}$, for fixed $r$ and either choice of $s$, so the Auslander algebra is generated by the set $\{\phi_{r}^{w}\}$ with $r\in\mathbb{Z}/2\mathbb{Z}$ and $w\in W$. However, for non-empty word, note each $\phi_{r}^{w}$ is generated as a product
$\phi_{r}^{w}=\phi_{r+p(|w|)}^{w_{1}}\cdots\circ\phi_{r+1}^{w_{|w|-1}}\circ\phi_{r}^{w_{|w|}}$
@@ 3.1.1. Full projective subcategory of an abelian category with enough projectives
For a finite tensor category $\mathcal{C}$, let $\mathcal{P}_{\mathcal{C}}$ denote its full projective subcategory. Since $\mathcal{P}_{\mathcal{C}}$ need not be unital nor abelian, we wish to consider a category with objects of this form. Thus, let $\mathcal{B}_{R}$ denote the category with objects additive tensor subcategories of finite tensor categories with maps right exact tensor functors up to natural equivalence. Let $\mathcal{A}$ denote the full subcategory of $\mathcal{B}_{R}$ consisting of finite tensor categories.
More generally, suppose $\mathcal{C}$ is an abelian category with enough projectives and possibly some additional structures, and let $\mathcal{P}_{\mathcal{C}}$ denote its full subcategory of projectives. Let $\mathcal{B}_{R}$ be the category with objects additive categories with the same structure and morphisms natural equivalence classes of right exact functors preserving this structure, and let $\mathcal{A}$ denote the full subcategory of $\mathcal{B}_{R}$ with abelian objects.
@@ 3.3. Categorical reconstruction from projectives
If $\mathcal{P}$ is a full projective subcategory of a finite tensor category, then for a nonzero map $\phi:P_{i}\to P_{k}$ between projective indecomposables, im $\text{im\,}\phi$ is a simple object if and only if for every map nonzero $\psi:P_{l}\to P_{k}$ and $j:P_{k}\to P$ such that $j\circ\psi=0$, then $j\circ\phi=0$ as well where $P_{i},P_{k},P_{l}$ are projective indecomposables and $P$ is projective.
@@ 4.1.1. The map $\pi_{*}:r(K_{2n})\to r(DK_{n})_{0}$
Thus, since $K-\overline{K}$ is in the annihilator of all simple tensors, it is in the annihilator of the entire tensor product. Thus, if $[M_{1}],[M_{2}]\in r(DK_{n})_{0}$, then $[M_{1}\otimes M_{2}]\in r(DK_{n})_{0}$. ∎
@@ 5.1.3. Indecomposable modules of $K_{2}$
We separate the indecomposable modules of $K_{2}$ into three types: the simple-projective $\{V(r),P(r)\}$, the Syzygy type $\{\Omega^{s}V(r),\Omega^{-s}V(r)\}$, and the continuous M-type $\{M_{n}(r,\eta)\}$. The dimensions of these modules are:
$\dim V(r)=1,\dim P(r)=4,\dim\Omega^{\pm s}V(r)=2s+1,\dim M_{n}(r,\eta)=2n.$
We deduce the following from [3].
@@
    $M_{s}(r,\eta)\otimes M_{n}(r^{\prime},\eta^{\prime})=\oplus snP(r+r^{\prime})$ if $\eta\neq\eta^{\prime}$, $M_{s}(r,\eta)\otimes M_{n}(r^{\prime},\eta)=\oplus s(n-1)P(r+r^{\prime})\oplus M_{s}(0,\eta)\oplus M_{s}(1,\eta)$ if $s\leq n$.
@@ 5.2. Tensor ideals of the Representation Category of $K_{2}$
Recall for any projective $P\in\operatorname{Rep}(K_{2})$ and indecomposable $A\in\operatorname{Rep}(K_{2})$, $A\otimes P$ is projective. Also, $P(0)$ and $P(1)$ are the projective modules of $K_{2}$ from Section 2.2.2. Further, note that by Propositions 2.4 ($V(r^{\prime})$), 2.25 and 2.33 $(M_{n}(r^{\prime},\eta^{\prime})$), Corollary 2.22 $(\Omega^{n}(V(r))$, Corollary 2.24 $(\Omega^{-n}(V(r))$ and Corollary 2.10 ($P(r^{\prime})$) of [3], for any indecomposable $A\in\operatorname{Rep}(K_{2})$,
@@
A tensor ideal $\mathcal{T}$ of $\operatorname{Rep}(K_{2})$ is fully characterized by the set $A_{\mathcal{T}}$ where $A_{\mathcal{T}}$ is a subset of the set of all finite subsets of +×$\mathbb{Z}_{+}\times\mathbb{C}P^{1}\times\mathbb{Z}_{+}\times\{0,1\}$ where $T\in A_{\mathcal{T}}$ if and only if $\sum\limits_{(n,\eta,k,r)\in T}kM_{n}(r,\eta)\in\mathcal{T}$.